A man can row three-quarters of a kilometer against the stream in 45/4 minutes. The speed (in km/hr) of the man in still water ?
Answer by Alan3354(69443)

A man can row three-quarters of a kilometer against the stream in 45/4 minutes. The speed (in km/hr) of the man in still water ?
----------
r = rowing speed (in still water)
c = current
-----------
r - c = d/t = (3/4)km/(45/4)/60))hr
r - c = (3/4)/(45/240) km/hr
r - c = 720/135 km/hr
r - c = 16/3 km/hr
r = c + 16/3 km/hr
-------------------
He can row 16/3 km/hr plus whatever c for the current is.
The current wasn't spec'ed.